OpenStudy (anonymous):

Determine if the Mean Value Theorem applies to the function on the given interval. If the either applies, find all values of c that are guaranteed by the theorem f(x)=x^3-x^2-2x on [-1,1]

OpenStudy (science0229):

First find f'(x), which is 3x^2-2x-2. Now, you have to find all values of c that satistify\[\frac{ f(1)-f(-1) }{ 1-(-1) }=\frac{ -2-0 }{ 2 }=-1=f'(c)=3c^2-2c-2\] Solve for c (-1<c<1).

OpenStudy (mathmale):

We need to ask ourselves, first, whether the function f(x) = 3x^2-2x-2 is continuous and differentiable on the given interval; those are the conditions on which the Mean Value Theorem is based.